OpenStudy (anonymous):

how do you do this kind of problems : one number is 3 more than half another number. the sum of the numbers is 15

OpenStudy (amistre64):

x + y = z x = 3y z = 15 3y + y = 15

OpenStudy (amistre64):

almost had it, gotta be good at reading

OpenStudy (amistre64):

x = y/2+3

OpenStudy (amistre64):

y/2 +3 +y = 15 y +6 +2y = 30 3y = 24 y = 8

OpenStudy (amistre64):

x = 7 then; since 8/2 = 4 + 3 = 7